A van, stopped at a traffic light, is hit directly in the rear by another car traveling at 4 m/s. Assume that the transmission of the van is in neutral, the brakes are not being applied, and the collision is elastic. If the mass of the car is 835 kg, and the final velocity of the car is -0.4 m/s, what is the mass of the van kg
and what is the final velocity of the van?
Answer #1

momentum conservation 835X4= (8358 -0.ult (mo mu= 3674 -0 By Energy conservation: - 2 m, u = 1 mart 1 mare? o mp=13226.4 @ te